OpenStudy (anonymous):

how do you solve this by taking the square root of both sides. Express irrational answers in radical form. (x-10)to the 2nd power=4

OpenStudy (anonymous):

\[(x-10)^2=4\implies x-10=\pm \sqrt{4}\implies x=10\pm \sqrt{4}\]

OpenStudy (amistre64):

well; by taking the sqrt of each side; the sqrt(yada^2) cancels each other to just "yada"

OpenStudy (amistre64):

but it tends to produce spurious results which should be double checked for accuracy